This monument in the shape of a square pyramid was built in a park.
The height of the pyramid is 20ft. The base length of the pyramid is 24ft.
The four lateral faces of the pyramid will be painted. Both paint types will be used on the monument.
The cost per gallon of the regular paint is $35. The Paint coverage (square feet per gallon) is 200 for the regular paint.
The cost per gallon of the metallic paint is $100. The Paint coverage (square feet per gallon) is 80 for the metallic paint.
The regular paint will require 2 coats, and the metallic paint will require 1 coat. Both paint types are sold only by the gallon. What is the least amount of money that should be budgeted to paint the monument?

Answer by mananth(16946)


Regular paint 2 coats 200sq.ft /gallon
1119.54*2/200= 11.2 gallons = 12 gallons
12*35= 390
1119.54/80= 13.99 gallons - 13 gallons
13*100 =1300

390+1300= $1690 Budget